Manufacturing cost data for Orlando Company, which uses a job order cost system, are presented below.

Indicate the missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct labor cost and the rate is the same. (Round overhead rate to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 5,275.)

Case A

Case B

Case C

Direct materials used $enter a dollar amount (a) $91,700 $69,550
Direct labor 56,000 148,500 enter a dollar amount (h)
Manufacturing overhead applied 44,800 enter a dollar amount (d) enter a dollar amount (i)
Total manufacturing costs 154,650 enter a dollar amount (e) 218,950
Work in process 1/1/17 enter a dollar amount (b) 19,200 18,900
Total cost of work in process 204,700 enter a dollar amount (f) enter a dollar amount (j)
Work in process 12/31/17 enter a dollar amount (c) 13,300 enter a dollar amount (k)
Cost of goods manufactured 194,800 enter a dollar amount (g) 231,400
Answer #1
Case A Case B Case C
Direct materials used 53,850 (a) 91,700 69,550
Direct labor 56,000 148,500 83000 (h)
Manufacturing overhead applied 44,800 118800 (d) 66400 (i)
Total manufacturing costs 154,650 359,000 (e) 218,950
Work in process 1/1/17 50,050 (b) 19,200 18,900
Total cost of work in process 204,700 378,200 (f) 237,850 (j)
Work in process 12/31/17 9,900 (c) 13,300 6,450 (k)
Cost of goods manufactured 194,800 364,900 (g) 231,400
Case A
(a) Direct materials used = Total manufacturing costs + Direct labor - Manufacturing overhead applied
= 154650-56000-44800
= 53850
(b) Work in process 1/1/17 = Total cost of work in process - Total manufacturing costs
= 204700 - 154650
= 50050
( c)Work in process 12/31/17 = Total cost of work in process - Cost of goods manufactured
= 204700 - 194800
= 9900
Case B
Overhead rate = Manufacturing overhead applied / Direct labor
= 44800/56000
= 0.8
(d) Manufacturing overhead applied = Direct labor * Overhead rate  
= 148500 * 0.8
= 118800
( e)Total manufacturing costs = Direct materials used + Direct labor + Manufacturing overhead applied
= 91700 + 148500 + 118800
= 359000
(f)Total cost of work in process = Total manufacturing costs + Work in process 1/1/17
= 359000+19200
= 378200
(g) Cost of goods manufactured = Total cost of work in process - Work in process 12/31/17
= 378200-13300
= 364900
Case C
Direct labor + Manufacturing overhead applied = Total manufacturing costs - Direct materials used
or, Total manufacturing costs - Direct materials used = 218950 - 69550
= 149400
(h) Direct labor = 149400*1/1.8
= 83000
(i) Direct materials used = 149400*0.80/1.80
= 66400
(j) Total cost of work in process = Total manufacturing costs + Work in process 1/1/17
= 218950+18900
= 237850
(k)Work in process 12/31/17 = Total cost of work in process - Cost of goods manufactured
= 237850-231400
= 6450